OpenStudy (anonymous):

Gabriel is comparing three investment accounts offering different rates. Account A: APR of 2.85% compounding monthly Account B: APR of 2.98% compounding quarterly Account C: APR of 2.90% compounding daily Which account will give Gabriel at least a 3% annual yield?

OpenStudy (anonymous):

account c is the best deal

OpenStudy (anonymous):

its the classic would u rather have a million dollars or a penny that doubles everyday scenario

jimthompson5910 (jim_thompson5910):

the answer isn't C, but you would use this formula APY = (1+r/n)^n - 1 where r = APR n = compounding frequency

OpenStudy (anonymous):

the anser is soo C

OpenStudy (anonymous):

it would make you the most money

OpenStudy (anonymous):

the question basically asks which is best since it says at least 3 percent

jimthompson5910 (jim_thompson5910):

APY = (1+r/n)^n - 1 APY = (1+0.029/365)^365 - 1 APY = 0.02942340858506 APY = 2.942% so it's not 3% or over 3%

OpenStudy (anonymous):

just think about which one has the best value

jimthompson5910 (jim_thompson5910):

well this is counter-intuitive